# Seek in the Dark: Reasoning via Test-Time Instance-Level Policy Gradient in Latent Space

### v2

@@ 3.3 Ideal Experiment: Perfect Sparse Reward Model
To further evaluate the effectiveness of exploration within the latent space, we conduct experiments using a Perfect Sparse Reward Model (PSRM). In this experimental setup, a reward value of 0 is assigned exclusively when the generated final answer exactly matches the ground truth. In all other cases, a reward of 1-1 is given. The results of these experiments are presented in Table 1. Despite the limited feedback provided by this sparse reward signal—particularly in contrast to the self-reward model—the PSRM demonstrates notable performance gains. Specifically, it achieves an average improvement of 19.12% score over the CoT method and surpasses the self-reward version by an average of 12.57% score. These results highlight the potential of LatentSeek, even when guided by sparse reward.
Notably, when Qwen2.5-1.5B-Instruct is employed as the backbone, the performance under the PSRM often matches or exceeds that of CoT using the larger Qwen2.5-7B-Instruct. For example, on GSM8K (Prompt 2), the former achieves a score of 67.48% compared to 66.41% for the latter. Similarly, on MATH-500 (Prompt 1), the smaller model reaches 75.80%, surpassing the larger model’s score of 72.80%. These findings suggest that smaller models have acquired substantial knowledge but may lack effective mechanisms to elicit them. Our proposed approach offers a mechanism for activating implicit knowledge, thereby enhancing the model’s ability.
@@ 3.4 Test-Time Scaling: scaling up the iteration of LatentSeek
In addition to increasing the number of generated tokens at test time Muennighoff et al. (2025), Snell et al. (2025), we propose an alternative approach to test-time scaling: increasing the number of LatentSeek iterations. The relationship between model performance on GSM8K Cobbe et al. (2021) and the number of iterations is illustrated in Figure˜2. As shown, the ideal reward model yields a consistently improving trend and outperforms the self-reward method across all model backbones, suggesting that test-time scaling can be achieved without necessitating a dense reward function in our setting. Conversely, the self-reward method exhibits rapid initial gains followed by a plateau, attributable to the limitations in the reward model’s accuracy. Notably, with an increased number of iterations, performance eventually surpasses this plateau and continues to improve. These findings indicate that, given an appropriate reward model, searching through the latent space offers a promising new direction for test-time scaling.
@@
Following the methodology proposed by Liu et al. (2025), we conduct extreme scaling experiments using the Qwen2.5-1.5B-Instruct model Team (2024a). In contrast to Liu et al. (2025), which leverages Process Reward Models (PRM), our approach is specifically tailored to align with Outcome Reward Models (ORM). Given the current limitations in ORM design (see Section˜E.5) and the inherent challenges of self-rewarding in small-scale models, we investigate the effectiveness of this novel test-time scaling strategy through the use of the PSRM. As shown in Table˜3, Qwen2.5-1.5B-Instruct achieves a 14% point performance gain over GPT-4o on the AIME2024 benchmark. Furthermore, on the MATH-500 dataset, it attains the highest overall accuracy among all evaluated models, with a marginal 2.7% point gap behind o1-preview. Remarkably, the model requires only an average of 61.8 iterations on MATH-500 to reach the performance, outperforming a 256 compute consumption of explicit space scaling methods in terms of efficiency. This implies that, under an appropriate reward model, the latent space represents a more efficient option for test-time scaling compared to the explicit space.
@@ 3.5 Algorithmic Statistics
Table˜4 reports two key statistical features: (1) the ratio of final answer tokens to the original CoT tokens, and (2) the average number of update iterations across the entire dataset. In nearly all cases, excluding Qwen2.5-1.5B-Instruct on GSM8K with prompt 2, the token ratio does not exceed 1.1. This suggests that, unlike long CoT approaches Xu et al. (2025b), our method does not rely on generating extended outputs to achieve strong performance, thus avoiding overly verbose generation. Moreover, concerning the number of update iterations, which constitutes the core computational component of the algorithm, the average iteration counts for GSM8K and MATH-500 across various model backbones and prompts are 0.86 and 1.23, respectively. Both values are below 2, indicating that the search process in the latent space typically converges in fewer than two iterations for average-level questions. This highlights the computational efficiency and rapid convergence behavior of our method.
@@ 3.6 Qualitative Analysis
To gain deeper insight into model-generated reasoning, we analyze the distribution of the first three tokens produced under Prompt 1 in the GSM8K dataset Cobbe et al. (2021) using the Llama3.1-8B-Instruct model. The frequent appearance of nonsensical or obscure tokens—such as “thecy,” “theella,” and “Sid”—indicates that the model’s reasoning mechanisms often diverge from human cognitive strategies, even when producing superficially coherent outputs. A representative example, detailed in Table˜5, further illustrates this divergence: despite generating linguistically anomalous expressions such as “total downloads of downloads,” and “down step two andLet,” the model still arrives at the correct answer. These findings suggest that the optimal reasoning trajectories for language models are fundamentally different from those of humans and may be more effectively navigated within the model’s latent space. For further discussion, see Appendix˜H.
Table 5: Case Studies: the questions, the answers, the CoT generations, and the LatentSeek generations are listed. The incoherence examples are marked with blue color. See Appendix˜H for more cases.
@@ 4 Related Work
Recent advances in reasoning capabilities of large language models (Brown et al., 2020, OpenAI, 2023) have been enhanced through prompting techniques. CoT prompting (Wei et al., 2022, Kojima et al., 2022, Zhou et al., 2022, Li et al., 2024) encourages models to generate intermediate reasoning steps. Unlike these static approaches, our method dynamically optimizes the reasoning process for each problem instance. Compute-optimal scaling (Snell et al., 2025, Misaki et al., 2025) adaptively adjusts inference strategies based on task complexity. Latent CoT methods (Hao et al., 2024, Shen et al., 2025, Cheng and Van Durme, 2024, Deng et al., 2024) replace explicit text-based reasoning with continuous representations. The broader field of learning to reason includes techniques like process supervision (Uesato et al., 2022) and self-critique (Huang et al., 2022).
@@
The integration of Reinforcement Learning and LLM starts from the realm of Human Feedback (RLHF) (Ouyang et al., 2022, Bai et al., 2022), which commonly employs algorithms such as Proximal Policy Optimization (PPO) (Schulman et al., 2017). Recent advancements include Direct Preference Optimization (Rafailov et al., 2023), Statistical Rejection Sampling (Liu et al., 2023b), and GRPO (Guo et al., 2025), which specifically address reasoning tasks. Innovations in reward modeling have explored human preference-based architectures (Schulman et al., 2017, Hazra et al., 2024, Yuan et al., 2022), automated language model-driven design (Kwon et al., 2023), and multi-agent verification frameworks (Lifshitz et al., 2025). Chen et al. (2024b) employs a variational method to improve the fitting of latent trajectories by updating model parameters. While these methods focus on modifying model parameters during training, our approach optimizes latent representations during testing, without altering the model parameters.
@@
Various approaches have been developed for controlling language model outputs, including conditioning on control codes (Keskar et al., 2019), gradient-based steering (Dathathri et al., 2019), and prompt optimization (Qin et al., 2023). At test time, techniques like self-consistency (Wang et al., 2022), recitation-augmented generation (Sun et al., 2023), and test-time alignment Zhang et al. (2025) improve outputs through sampling and selection. Some Test-Time Training (TTT) framewroks (Sun et al., 2020; 2024, Hardt and Sun, 2023) integrate self-supervised objectives for online model updates. Our work differs from test-time planning approaches (Hao et al., 2023) by optimizing in a continuous latent space rather than performing a discrete search.
@@
Prompt Tuning and Prefix Tuning adapt language models by prepending trainable vectors to inputs or hidden states Lester et al. (2021), Liu et al. (2024a), Li and Liang (2021), Liu et al. (2021). However, both require labeled data and full backpropagation, incurring high computational cost. In contrast, our method leverages latent-space manipulation without training data or model updates, enabling efficient, flexible adaptation.
@@ C.2 Theoretical Analysis: Independent Updating
  • LatentSeek: the optimized tokens, as they are updated independently, act like the provers who jointly determine the outcome.
  • MIP-Bounded: Different from MIP, each prover in LatentSeek can only generate 1 token, whose number of bits is bounded. Furthermore, a polynomial proportion of tokens is allowed to be optimized, leading to a polynomial number of provers.
@@ E.1 LatentSeek vs BoN (N=10)
In Table˜1, we show that Best-of-N (BoN) is outperformed by LatentSeek, even though BoN consumes a comparable compute consumption (See Figure˜4). One might also ask how LatentSeek compares to BoN when $N=10$; in this setting, BoN requires at least five times more sequence-level computation than LatentSeek. As reported in Table˜6, LatentSeek remains superior on all datasets and all backbone models.
@@ E.2 Qwen3 AIME Performance
This section provides a detailed study on the performance of the Qwen3-4B-Instruct-2507 model on the challenging AIME2024 dataset. Figure˜3 illustrates the accuracy of LatentSeek compared against the Chain-of-Thought (CoT) and Best-of-N (BoN) baselines under two prompt formats.
@@ E.3 Token Efficiency Comparison with BoN Baseline
As shown in Figure˜4, LatentSeek achieves higher token efficiency than BoN in all of cases across different model families and sizes, while attaining superior reasoning accuracy as reported in Table˜1. These results demonstrates that LatentSeek can enhance performance without incurring additional token overhead for all models, highlighting the efficiency of latent space optimization.
@@
Figure 4: Comparison of average tokens consumed per problem between LATENTSEEK and the Best-of-N (BoN) baseline when using Prompt 2. The token consumption is comparable for the two methods, with BoN consuming slightly more tokens in total. For experimental results where BoN consumes extremely more calculation, please refer to Table˜6.
@@ E.4 Greedy Decoding vs Sampling
Our experiments show that greedy decoding closely approximates categorical sampling while being a better choice for reducing stochastic error. As detailed in Table˜7, greedy decoding performs slightly better than the sampling method. We hypothesize that this is because sampling can amplify the approximation error when replacing an expected mean with an empirical one. Therefore, we use greedy decoding for our main experiments.

### v3

@@ 3.3 Ideal Experiment: Perfect Sparse Reward Model
Table 3: Accuracy results with Perfect Sharp Reward Model (PSRM). Self: the self-reward mechanism. The reported score is the average of both prompts.
Reward Qwen2 Qwen2.5 LLaMA3.1 Avg
7B 1.5B 7B 14B 8B
GSM8K
LatentSeek-Self 82.3 57.5 87.6 92.6 81.1 80.2
LatentSeek-Perfect Sharp Reward Model 92.8+10.5 76.5+19.0 93.9+6.3 96.0+3.4 91.7+10.6 90.2+10.0
MATH-500
LatentSeek-Self 51.1 45.0 66.6 75.5 50.8 57.8
LatentSeek-Perfect Sharp Reward Model 71.8+20.7 66.7+21.7 81.8+15.2 86.6+11.1 70.5+19.7 75.5+17.7
AIME2024
LatentSeek-Self 3.3 5.0 13.3 13.3 8.3 8.6
LatentSeek-Perfect Sharp Reward Model 8.3+5.0 6.7+1.7 15.0+1.7 25.0+11.7 10.0+1.7 13.0+4.4
To inspect the value of exploration in latent space, we evaluate a Perfect Sharp Reward Model (PSRM) that returns 0 only when the final answer exactly matches the ground truth and 1-1 otherwise. This yields an all-or-nothing signal: until the correct answer is reached, every step receives identical feedback. Results (Table˜3) show that, despite this extreme sharpness, the PSRM still delivers substantial improvements, outperforming the self-reward variant by an average of 10.67 points across all evaluated settings. Conceptually, this setup reduces optimization to maximizing the success rate under a near-constant reward landscape; the search carries almost no directional information and is therefore close to blind exploration. This demonstrate that exploration in latent space alone can yield meaningful performance improvements, even under sharp feedback.
Small Model and Large Iterations.
Following Liu et al. (2025), we analyze the small Qwen2.5-1.5B-Instruct (Team, 2025) under a large iteration budget to probe the limits of latent-space exploration. As shown in Table˜4, Qwen2.5-1.5B-Instruct yields a 14-point gain over GPT-4o on AIME2024. On MATH-500, it achieves a high accuracy score, trailing o1-preview by only 2.7 points. These results demonstrate that even a 1.5B-parameter model’s latent space is highly expressive and that LatentSeek can be an effective mechanism for eliciting this expressivity.
@@ 3.4 Test-Time Scaling: scaling up the iteration of LatentSeek
In addition to increasing the number of generated tokens at test time Muennighoff et al. (2025), Snell et al. (2025), we propose an alternative approach to test-time scaling: increasing the number of LatentSeek iterations. The relationship between model performance on GSM8K Cobbe et al. (2021) and the number of iterations is illustrated in Figure˜2. As shown, the ideal reward model yields a consistently improving trend and outperforms the self-reward method across all model backbones, suggesting that test-time scaling can be achieved without necessitating a dense reward function in our setting. Conversely, the self-reward method exhibits rapid initial gains followed by a plateau, attributable to the limitations in the reward model’s accuracy.Notably, as the number of iterations increases further, performance eventually surpasses this plateau and continues to improve. The above shows that test-time scaling remains attainable even under a sharply peaked reward. It further suggests that, given an appropriate reward model, searching through the latent space offers a promising new direction for test-time scaling.
@@
Following the methodology proposed by Liu et al. (2025), we conduct extreme scaling experiments using the Qwen2.5-1.5B-Instruct model Team (2025). In contrast to Liu et al. (2025), which leverages Process Reward Models (PRM), our approach is specifically tailored to align with Outcome Reward Models (ORM). Given the current limitations in ORM design (see Section˜E.5) and the inherent challenges of self-rewarding in small-scale models, we investigate the effectiveness of this novel test-time scaling strategy through the use of the PSRM. As shown in Table˜4, Qwen2.5-1.5B-Instruct achieves a 14% point performance gain over GPT-4o on the AIME2024 benchmark. Furthermore, on the MATH-500 dataset, it attains the highest overall accuracy among all evaluated models, with a marginal 2.7% point gap behind o1-preview. Remarkably, the model requires only an average of 61.8 iterations on MATH-500 to reach the performance, outperforming a 256 compute consumption of explicit space scaling methods in terms of efficiency. This implies that, under an appropriate reward model, the latent space represents a more efficient option for test-time scaling compared to the explicit space.
3.5 Comparison with Unguided Search
Table 5: Comparison with Unguided Search on GSM8K. SE: Stochastic Exploration.
Model Qwen2 Qwen2.5 LLaMA3.1 Avg
7B 1.5B 7B 14B 8B
GSM8K: Prompt 1
SE ($\sigma^{2}=0.50$) 66.19 64.22 88.60 92.33 49.73 72.21
SE ($\sigma^{2}=0.75$) 65.96 63.53 88.60 92.03 49.73 71.97
SE ($\sigma^{2}=1.00$) 65.13 63.76 89.08 91.88 50.57 72.08
LatentSeek 84.38 70.89 90.14 92.49 78.54 83.29
GSM8K: Prompt 2
SE ($\sigma^{2}=0.50$) 64.32 16.68 65.60 92.04 67.85 61.30
SE ($\sigma^{2}=0.75$) 64.77 15.69 64.92 91.89 67.02 60.86
SE ($\sigma^{2}=1.00$) 63.84 16.07 66.34 92.12 66.72 61.02
LatentSeek 80.21 44.20 85.06 92.72 83.70 77.18
We compare LatentSeek with a single-step stochastic explorative search in latent space on the GSM8K (Cobbe et al., 2021), to assess the effectiveness of our gradient-guided search. In this baseline, instead of performing iterative gradient ascent, a Gaussian noise is added once to the initial latent representations,
$z\leftarrow z+\epsilon,\,\epsilon\sim\mathcal{N}(0,\sigma^{2}I)$
, and evaluate three noise levels $\sigma^{2}=0.5,0.75,$ and $1.0$. As shown in Table˜5, LatentSeek outperforms the stochastic exploration baselines by an average of 13.66 points, indicating that the gains arise from effective gradient guidance rather than random perturbations.
3.6 Analysis of Self-Reward Mechanism
Table 6: Self-Reward Mechanism Percentage Accuracy on GSM8K.
Backbone Qwen2 Qwen2.5 LLaMA3.1 Avg
7B 1.5B 7B 14B 8B
Acc 85.78 77.15 80.17 89.83 80.46 82.68
[…]
Table 7: Comparison with Constant Reward baseline
Backbone Qwen2 Qwen2.5 LLaMA3.1 Avg
7B 1.5B 7B 14B 8B
LatentSeek (Constant) 60.19 29.49 64.52 90.67 69.98 62.97
LatentSeek (Self) 80.21 44.20 85.06 92.72 83.70 77.18
$\Delta$ +20.02 +14.71 +20.54 +2.05 +13.72 +14.21
[…]
@@ 3.7 Algorithmic Statistics
Table˜8 reports two key statistical features: (1) the ratio of final answer tokens to the original CoT tokens, and (2) the average number of update iterations across the entire dataset. In nearly all cases, excluding Qwen2.5-1.5B-Instruct on GSM8K with prompt 2, the token ratio does not exceed 1.1. This suggests that, unlike long CoT approaches Xu et al. (2025b), our method does not rely on generating extended outputs to achieve strong performance, thus avoiding overly verbose generation. Moreover, concerning the number of update iterations, which constitutes the core computational component of the algorithm, the average iteration counts for GSM8K and MATH-500 across various model backbones and prompts are 0.86 and 1.23, respectively. Both values are below 2, indicating that the search process in the latent space typically converges in fewer than two iterations for average-level questions. This highlights the computational efficiency and rapid convergence behavior of our method. For a FLOPs analysis of our method, please refer to Appendix˜G for details.
@@ 3.8 Qualitative Analysis
To gain deeper insight into model-generated reasoning, we analyze the distribution of the first three tokens produced under Prompt 1 in the GSM8K dataset Cobbe et al. (2021) using the Llama3.1-8B-Instruct model. The frequent appearance of nonsensical or obscure tokens—such as “thecy,” “theella,” and “Sid”—indicates that the model’s reasoning mechanisms often diverge from human cognitive strategies, even when producing superficially coherent outputs. A representative example, detailed in Table˜9, further illustrates this divergence: despite generating linguistically anomalous expressions such as “total downloads of downloads,” and “down step two andLet,” the model still arrives at the correct answer. These findings suggest that the optimal reasoning trajectories for language models are fundamentally different from those of humans and may be more effectively navigated within the model’s latent space. For further discussion, see Appendix˜I.
Table 9: Case Studies — the questions, the answers, the CoT generations, and the LatentSeek generations are listed. The incoherence examples are marked with blue color. See Appendix˜I for more cases.
@@
Figure 3: Examples of latent-space optimization trajectories. Blue: initial latent vector; Black: intermediate steps; Green: final latent vector; Red: optimization trajectory. For more trajectories, please refer to Figure˜11 and Figure˜12
To better understand the optimization dynamics within the continuous latent space, we conduct a visualization study of latent-vector trajectories. Specifically, for each token position $t$ and optimization step $k$, we collect the corresponding latent representations $z^{(k)}_{t}$ and apply Principal Component Analysis (PCA) to project these high-dimensional vectors into a three-dimensional space. The Figure˜9 highlights three representative patterns of latent-space optimization. The Blue dot denotes the initial latent vector, the Black dots indicate intermediate latent states, the Green dot marks the final latent vector, and the Red line traces the optimization trajectory. As illustrated in the figures, some token-specific latents exhibit zig-zag search, others follow a consistent unidirectional trajectory, and still others trace an V-shaped search path. (For more visualization, please refer to Figure˜11 and Figure˜12 in Appendix˜H)
@@ 4 Related Work
Recent advances in reasoning capabilities of large language models (Brown et al., 2020, OpenAI, 2023) have been enhanced through prompting techniques. CoT prompting (Wei et al., 2022, Kojima et al., 2022, Zhou et al., 2023, Li et al., 2024) encourages models to generate intermediate reasoning steps. Unlike these static approaches, our method dynamically optimizes the reasoning process for each problem instance. Compute-optimal scaling (Snell et al., 2025, Inoue et al., 2025) adaptively adjusts inference strategies based on task complexity. Latent CoT methods (Hao et al., 2024, Shen et al., 2025, Cheng and Van Durme, 2024, Deng et al., 2024) replace explicit text-based reasoning with continuous representations. The broader field of learning to reason includes techniques like process supervision (Uesato et al., 2022) and self-critique (Yang et al., 2024b).
@@
The integration of Reinforcement Learning and LLM starts from the realm of Human Feedback (RLHF) (Ouyang et al., 2022, Bai et al., 2022), which commonly employs algorithms such as Proximal Policy Optimization (PPO) (Schulman et al., 2017). Recent advancements include Direct Preference Optimization (Rafailov et al., 2023), Statistical Rejection Sampling (Liu et al., 2024c), and GRPO (Guo et al., 2025), which specifically address reasoning tasks. Innovations in reward modeling have explored human preference-based architectures (Schulman et al., 2017, Hazra et al., 2025, Yuan et al., 2022), automated language model-driven design (Kwon et al., 2023), and multi-agent verification frameworks (Lifshitz et al., 2025). Chen et al. (2024a) employs a variational method to improve the fitting of latent trajectories by updating model parameters. While these methods focus on modifying model parameters during training, our approach optimizes latent representations during testing, without altering the model parameters.
@@
Various approaches have been developed for controlling language model outputs, including conditioning on control codes (Keskar et al., 2019), gradient-based steering (Dathathri et al., 2020). At test time, techniques like self-consistency (Wang et al., 2023), recitation-augmented generation (Sun et al., 2023), and test-time alignment Zhang et al. (2025) improve outputs through sampling and selection. Some Test-Time Training (TTT) framewroks (Sun et al., 2020; 2025, Hardt and Sun, 2024) integrate self-supervised objectives for online model updates. Our work differs from test-time planning approaches (Hao et al., 2023) by optimizing in a continuous latent space rather than performing a discrete search.
@@
Prompt Tuning and Prefix Tuning adapt language models by prepending trainable vectors to inputs or hidden states Lester et al. (2021), Liu et al. (2024d), Li and Liang (2021), Liu et al. (2022). However, both require labeled data and full backpropagation, incurring high computational cost. In contrast, our method leverages latent-space manipulation without training data or model updates, enabling efficient, flexible adaptation.
@@ C.2 Theoretical Analysis: Independent Updating
  • LatentSeek: the optimized tokens, as they are updated independently, act like the non-communicated provers who jointly determine the outcome. After decoding the latent vectors, where each prover proposes its own output, the subsequent autoregressive generation is regarded as part of the verifier function $V$. Regarding the CoT initialization, recall that the MIP framework permits communication among provers before the proving process begins. Accordingly, initializing the latent vectors with a CoT prompt can be interpreted as this pre-proof communication stage.
  • MIP-Bounded (LatentSeek): Different from MIP, each prover in LatentSeek can only generate 1 token, whose number of bits is bounded. Furthermore, a polynomial proportion of tokens (such as, $\rho=0.2$) is allowed to be optimized, leading to a polynomial number of provers.
@@
The collollary indicates that if the verifier and the provers satisfy some condition, the LatentSeek framework is strong enough, and the independence assumption can be amended. Intuitively speaking, the independence between the provers can be controlled by the verifier, such as the verifier can ask for dependence between the generated tokens.
There is a limitation in this theoretical analysis, which is why we present it in the appendix: it should be viewed only as an indication or heuristic guideline rather than a rigorous validation. In particular, we do not account for the change through gradient updates. As a result, there is no guarantee that every token becomes reachable in each step, especially when each prover relies solely on the gradients from the language-model head for guidance.
All in all, the theorem indicates that with a “strong” base model and a “well-aligned” reward model, the deficit introduced by the independence assumption can be effectively mitigated.
@@ E.1 LatentSeek vs BoN (N=10)
In Table˜1, we show that Best-of-N (BoN) is outperformed by LatentSeek, even though BoN consumes a comparable compute consumption (See Figure˜5). One might also ask how LatentSeek compares to BoN when $N=10$; in this setting, BoN requires at least five times more sequence-level computation than LatentSeek. As reported in Table˜10, LatentSeek remains superior on all datasets and all backbone models.
@@ E.2 Qwen3 AIME Performance
This section provides a detailed study on the performance of the Qwen3-4B-Instruct-2507 model on the challenging AIME2024 dataset. Figure˜4 illustrates the accuracy of LatentSeek compared against the Chain-of-Thought (CoT) and Best-of-N (BoN) baselines under two prompt formats.
@@ E.3 Token Efficiency Comparison with BoN Baseline
As shown in Figure˜5, LatentSeek achieves higher token efficiency than BoN in all of cases across different model families and sizes, while attaining superior reasoning accuracy as reported in Table˜1. These results demonstrates that LatentSeek can enhance performance without incurring additional token overhead for all models, highlighting the efficiency of latent space optimization.
@@
Figure 5: Comparison of average tokens consumed per problem between LATENTSEEK and the Best-of-N (BoN) baseline when using Prompt 2. The token consumption is comparable for the two methods, with BoN consuming slightly more tokens in total. For experimental results where BoN consumes extremely more calculation, please refer to Table˜10.
@@ E.4 Greedy Decoding vs Sampling
Our experiments show that greedy decoding closely approximates categorical sampling while being a better choice for reducing stochastic error. As detailed in Table˜11, greedy decoding performs slightly better than the sampling method. We hypothesize that this is because sampling can amplify the approximation error when replacing an expected mean with an empirical one. Therefore, we use greedy decoding for our main experiments.